Bass - Samples: The bass in Clear Lake have weights that are normally distributed with a mean of 2.2 pounds and a standard deviation of 0.8 pounds. Suppose you catch a stringer of 6 bass with a total weight of 16.5 pounds. Here we determine how unusual this is.

(a) What is the mean fish weight of your catch of 6? Round your answer to 1 decimal place.

(b) If 6 bass are randomly selected from Clear Lake, find the probability that the mean weight is greater than the mean of those you caught. Round your answer to 4 decimal places.
Answer #1

a) Mean weight of 6 fishes = (Total weight)/6 = 16.5/6 = 2.75 = 2.8

b) P(Mean > 2.75)

= P(z > 1.68)

= 0.0461